OpenStudy (anonymous):

SOLVE FOR T: 6P=Pe^(1.4t)

Nnesha (nnesha):

first we need to get rid of p at right side

Nnesha (nnesha):

how would you cancel out p from right side ?

Nnesha (nnesha):

do you know ? yes or no ?

OpenStudy (anonymous):

No, I don't know

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

it's p `times` e^P(1.4t) so to cancel out p from right side do opposite of `multiplication `

OpenStudy (anonymous):

Divide by P on both sides to cancel P?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

\[\huge\rm \frac{ 6p }{ p}=\frac{ pe^{(1.4t)} }{ p }\] what do you get ?

OpenStudy (anonymous):

6=e^(1.4t)

Nnesha (nnesha):

yes right not take natural log both sides to cancel out the e \[\huge\rm \cancel{ln e}^a = a\]bec ln and e cancel each other out

OpenStudy (anonymous):

t= 1.2798 ?

Nnesha (nnesha):

hm no how did you get that

Nnesha (nnesha):

oh wait

OpenStudy (anonymous):

ln of 6 divided by 1.4? is that wrong?

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

Yay!!!

Nnesha (nnesha):

Yahoo!!! good job!

OpenStudy (anonymous):

thanks so much!

Nnesha (nnesha):

yw :=)